# Visibly Fair Mechanisms

### v1

@@ 6 Efficiency
Consider two officers $I=\{i_{1},i_{2}\}$ of the same type $t$, and two states $S=\{s_{1},s_{2}\}$ with capacities $q_{s_{1}}=2$ and $q_{s_{2}}=1$. Let there be a modular upper-bound of one officer of type $t$ allowed in $s_{1}$. Thus, the modular-induced message space partitions states into two zones: $z_{1}=\{s_{1}\}$ and $z_{2}=\{s_{2}\}$. Let the exogenous ranking of zones be $z_{1}\blacktriangleright_{i}z_{2}$ for both officers.
@@
    $a$ is feasible and respects the upper–bounds, and
@@
    there is no other allocation $a^{\prime}$ that is feasible and respects the modular upper–bounds such that for every officer $i$ with $a^{\prime}_{i}\neq a_{i}$ we have $a^{\prime}_{i}\succ_{i}a_{i}$.
A mechanism is constrained Pareto efficient if, for any problem, when agents submit truthful messages the outcome is constrained Pareto efficient.
@@
There might be no static mechanism that is visibly fair and constrained Pareto efficient.
Theorem 7 underscores a key tension in designing visibly fair mechanisms that must also adhere to constraints such as modular upper-bounds. At its core, this tension arises from the message-space design: enforcing both fairness and quota requirements demands that the space be curated to preempt scenarios where certain preference reports would necessarily violate upper-bounds. Consequently, to avoid such violations, the mechanism must collect more restricted preference information than might otherwise be desirable. This reduction in elicited information, while upholding fairness and preserving the bounds, can lead to efficiency losses because the mechanism may lack the information needed to detect and implement mutually beneficial reallocations that remain compliant with all constraints.
@@
Sequential Assignment: Process the officers in the order $(i_{1},\dots,i_{n})$ where $\pi(i_{1})<\cdots<\pi(i_{n})$. For each $k=1,\dots,n$:
@@
    $z_{1}\;=\;\left\{s\in S:\left(\forall\,(S_{h}^{t_{i_{k}}},k_{h}^{t_{i_{k}}})\in B_{k}^{t_{i_{k}}},\;s\notin S_{h}^{t_{i_{k}}}\right)\quad\text{and}\quad\left|\{j<k:a_{j}=s\}\right|<q_{s}\right\}$
@@
    Elicit from officer $i_{k}$ a message $m_{k}\in M_{i_{k}}$.
@@
    Let $a_{i_{k}}$ be officer $i_{k}$’s most-preferred state $s^{\ast}\in z_{1}$, according to $m_{i_{k}}$. Notice that this can be $\varnothing$ if $z_{1}=\varnothing$.
@@
    Update Binding Upper-Bounds:
    • Let $B^{t}_{k+1}=B^{t}_{k}$
    • For each upper-bound $h=(S_{h}^{t_{i_{k}}},k_{h}^{t_{i_{k}}})$ involving type $t_{i_{k}}$:
      • If $s^{\ast}\in S_{h}^{t_{i_{k}}}$:
          Calculate:
          $N_{h}=\left|\left\{i\in I:a^{k}_{i}\in S_{h}^{t_{i_{k}}},\ t_{i}=t_{i_{k}}\right\}\right|.$
          If $N_{h}=k_{h}^{t_{i_{k}}}$:
            Add $(S_{h}^{t_{i_{k}}},k_{h}^{t_{i_{k}}})$ to $B^{t}_{k+1}$
  5. 5.
    Proceed to the next officer $i_{k+1}$.
Outcome: After processing all officers $i_{1},\ldots,i_{n}$, the assignment $a^{n}=(a^{n}_{i})_{i\in I}$ is the mechanism’s final allocation:
$\psi\bigl{(}m\bigr{)}\;=\;a^{n}.$
The Dynamic Modular Priority Mechanism, therefore, sequentially assigns officers, dynamically tailoring the menu of available states for each officer based on previous assignments. At each officer’s turn, the mechanism constructs a zone containing all states with remaining capacity that do not belong to any subset of states where a modular upper-bound for the officer’s type has already become binding. The officer is then matched to her most-preferred state among these feasible alternatives. By construction, this ensures visible fairness since each officer always obtains her top choice from that zone. Furthermore, by dynamically adjusting the zones and enabling officers to fully express their preferences within these constraints, the mechanism attains constrained efficiency: any allocation that improves an officer’s assignment without harming others would necessarily violate at least one modular upper-bound. The following proposition summarizes these properties, including incentives. [^12] While the definitions of visible fairness and respecting modular upper bounds carry over directly without introducing additional notation, strategy-proofness warrants further clarification in our dynamic setting. Because the Dynamic Modular Priority Mechanism allows officers’ strategies to depend on both past history and the structure of the message spaces (i.e., the composition of the zones), our original, static definition of strategy-proofness does not explicitly cover this complexity. Nevertheless, the informal argument used in the proof of Theorem 8 presented in Appendix A remains sound: no matter how other players behave over time, an officer cannot secure a better outcome by misreporting, so truth-telling continues to be a dominant strategy.
@@ 7 Conclusion
We explored how designing mechanisms that restrict the preferences participants can report can help reconcile multiple policy objectives—particularly distributional constraints—with fairness principles grounded in strict priority orders. We introduced the concept of visible fairness, in which a mechanism never produces an allocation that appears to violate a participant’s priority based on the partially observed preferences. Our main theoretical result (Theorem 1) demonstrated that every visibly fair mechanism operates as a variant of serial dictatorship adapted to partial preference orders. We further showed that this framework can accommodate diverse distributional objectives by employing modular upper-bounds, which limit how many individuals of certain types can be placed in specified subsets of positions. Central to this approach are modular-induced message spaces, which prevent participants from specifying cross-group preference comparisons that would otherwise undermine these quotas.
We then characterized how these mechanisms can be made incentive-compatible by parsing out two critical conditions—expressiveness and (weak) availability—that together ensure no participant can profit by misreporting her partial preferences. While restricting the scope of preference reporting can, in principle, worsen efficiency, we introduced a dynamic modular-priority framework that sequentially elicits partial preferences and updates feasibility constraints in real time, thus recovering the best possible Pareto outcomes subject to capacity and distributional limits. Overall, this paper highlights how deliberately constraining participants’ message spaces provides mechanism designers with new levers for achieving desired policy goals while preserving fairness and incentive properties in priority-based allocation settings
@@ A.1 Proof of Theorem 1
Part 1: A $m$-queue allocation mechanism is visibly fair.
For any officer $k\in I$, since $a_{k}\in G(S^{k},m_{k})$, there does not exist a $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$. Meanwhile, for all $s\in S\setminus S^{k}$, |$|\{i\in I:i<k\text{ and }a_{i}=s\}|=q_{s}$. Therefore, the outcome of $m$-queue allocation mechanism is not visibly unfair.
@@
For any $m\in M$, consider the outcome $\psi(m)=a$ of a visibly fair mechanism. Define set $S^{k}$ for each officer $k\in I$ as follows:
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): Set $s^{k}=a_{k}$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
Since allocation $a^{k}$ is visibly fair, we also have that $a^{k}\in G(S^{k},m_{k})$. Since, $a_{k}\in G(S^{k},m_{k})$, $\psi$ is a $m$-queue allocation mechanism. ∎
@@ A.2 Proof of Theorem 2
For any officer $k\in I$, since $a_{k}\in G(S^{k},m_{k})$, there does not exist a $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$. Meanwhile, for all $s\in S\setminus S^{k}$, |$\left|\{i\in I:i<k\text{ and }a_{i}=s\}\right|=q_{s}$. Therefore, the outcome of partitioned priority mechanism cannot be visibly unfair.
@@
For each $m\in M$, consider the outcome $\psi(m)=a$ of a visibly fair mechanism. We now construct a zone selection function $\mathcal{C}_{k}$ and set $S^{k}$ for each officer $k\in I$ by using the following steps:
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $\mathcal{C}_{k}(S^{k},m)=z$ if $a_{k}\in z$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
The values of $\mathcal{C}_{k}(S^{\prime},m)$ for the remaining combinations of $S^{\prime}\subseteq S$ and $m\in M$ can take any value, since they are never used.
Since the allocation $a$ is visibly fair, we also have that for all $k$, $a^{k}\in G(S^{k},m_{k})$. Since, $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))$, $\psi$ is a partitioned priority mechanism.
@@ A.3 Proof of Theorem 3
For any officer $k\in I$, since $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m)$, there does not exist a $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$. Meanwhile, for all $s\in S\setminus S^{k}$, |$\left|\{i\in I:i<k\text{ and }a_{i}=s\}\right|=q_{s}$. Therefore, the outcome of ranked partitioned priority mechanism cannot be visibly unfair.
@@
Theorem 2’s constructed zone selection function and argument apply here as well. For each $m\in M$, consider the outcome $\psi(m)=a$ of a visibly fair mechanism. We now construct a zone selection function $\mathcal{C}_{k}$ and set $S^{k}$ for each officer $k\in I$ by using the following steps:
  • Step 00: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): $\mathcal{C}_{k}(S^{k},m)=z$ if $a_{k}\in z$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is |$\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
The values of $\mathcal{C}_{k}(S^{\prime},m)$ for the remaining combinations of $S^{\prime}\subseteq S$ and $m\in M$ can take any value, since they are never used.
Since the allocation $a$ is visibly fair, we also have that for all $k$, $a^{k}\in G(S^{k},m_{k})$. Since, $a_{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m))$, $\psi$ is a partitioned priority mechanism.
We need to additionally show that the constructed zone selection function is a ranked zone selection function.
Suppose not, then for officer $k\in I$, $S^{k}\cap\mathcal{C}_{k}(S^{k},m)=\left\{min\left(\mathcal{C}_{k}(S^{k},m),m_{k}\right)\right\}$ and there is a zone $z\in Z$ such that $z\rhd_{k}\mathcal{C}_{k}(S^{k},m)$ and $max\left(m_{k},z\right)\in S^{k}$. Since $max\left(m_{k},z\right)\succ_{m_{k}}min\left(\mathcal{C}_{k}(S^{k},m),m_{k}\right)$, $a^{k}=min\left(\mathcal{C}_{k}(S^{k},m_{k}),m_{k}\right)$ is visibly unfair for officer $k$ under $m$. ∎

### v2

@@ 6 Efficiency
Consider two officers $I=\{i_{1},i_{2}\}$ with priority $\pi(i_{1})<\pi(i_{2})$ and of the same type $t$, and two states $S=\{s_{1},s_{2}\}$ with capacities $q_{s_{1}}=2$ and $q_{s_{2}}=1$. Let there be a modular upper-bound of one officer of type $t$ allowed in $s_{1}$. Thus, the modular-induced message space partitions states into two zones: $z_{1}=\{s_{1}\}$ and $z_{2}=\{s_{2}\}$. Let the exogenous ranking of zones be $z_{1}\blacktriangleright_{i}z_{2}$ for both officers.
@@
    $a$ respects the upper–bounds, and
@@
    there is no other allocation $a^{\prime}$ that respects the modular upper–bounds such that for every officer $i$ we have $a^{\prime}_{i}\succsim_{i}a_{i}$.
A mechanism is constrained Pareto efficient if, for any problem, when agents submit any truthful message, the outcome is constrained Pareto efficient.
@@
Not all modular upper-bound constraints admit a (static) mechanism that is simultaneously visibly fair and constrained Pareto efficient.
The negative result in Theorem 7 is not, of course, universal. Standard SD satisfies those properties for modular upper-bound constraints that never bind. On the other hand, constant mechanisms can trivially respect constraints that result in a single feasible allocation which, by its uniqueness, would be constrained Pareto efficient.
Theorem 7, however, underscores a key tension in designing visibly fair mechanisms that must also adhere to non-trivial constraints such as modular upper-bounds. At its core, this tension arises from the message-space design: enforcing both fairness and quota requirements demands that the space be curated to preempt scenarios where certain preference reports would necessarily violate upper-bounds. Consequently, to avoid such violations, the mechanism must collect more restricted preference information than might otherwise be desirable. This reduction in elicited information, while upholding fairness and preserving the bounds, can lead to efficiency losses because the mechanism may lack the information needed to detect and implement mutually beneficial reallocations that remain compliant with all constraints.
@@
Quotas update procedure:
We next describe the procedure that updates the set of binding upper-bounds:
  • Let $\left(B^{t}_{k}\right)_{t\in T}$ be the current set of binding modular upper-bounds.
  • For each $(\Xi_{h},S_{h},k_{h})\in H$:
    • Let $N_{h}=\left|\left\{i\in I:a_{i}\in S_{h}\text{ and }t_{i}\in\Xi_{h}\right\}\right|$.
    • If $N_{h}=k_{h}$, then for every $t\in\Xi_{h}$, add $h$ to $B^{t}_{k}$.
Sequential Assignment: Process the officers in the order $(i_{1},\dots,i_{n})$. For each $k=1,\dots,n$:
@@
    $z_{1}\;=\;\left\{s\in S:\left(\forall\,(\Xi_{h},S_{h},k_{h})\in B_{k}^{t_{i_{k}}},\;s\notin S_{h}\right)\quad\text{and}\quad\left|\{j<k:a_{j}=s\}\right|<q_{s}\right\}$
@@
    Elicit from officer $i_{k}$ a message $m_{k}\in M_{i_{k}}$. Let $s^{\ast}$ be officer $i_{k}$’s most-preferred state in zone $z_{1}$, according to $m_{i_{k}}$.
@@
    Set $a_{i_{k}}=s^{\ast}$.
@@
    Follow the quotas update procedure and proceed to the next officer $i_{k+1}$.
Outcome: After processing all officers $i_{1},\ldots,i_{n}$, the assignment $a=(a_{i})_{i\in I}$ is the mechanism’s final allocation.
The Dynamic Modular Priority Mechanism, therefore, sequentially assigns officers, dynamically tailoring the menu of available states for each officer based on previous assignments. At each officer’s turn, the mechanism constructs a zone containing all states with remaining capacity that do not belong to any subset of states where a modular upper-bound for the officer’s type has already become binding. The officer is then matched to her most-preferred state among these feasible alternatives. By construction, this ensures visible fairness since each officer always obtains their top choice from that zone. Furthermore, by dynamically adjusting the zones and enabling officers to fully express their preferences within these constraints, the mechanism attains constrained efficiency: any allocation that improves an officer’s assignment without harming others would necessarily violate at least one modular upper-bound. The following proposition summarizes these properties, including incentives. [^19] While the definitions of visible fairness and respecting modular upper bounds carry over directly without introducing additional notation, strategy-proofness warrants further clarification in our dynamic setting. Because the Dynamic Modular Priority Mechanism allows officers’ strategies to depend on both history and the structure of the message spaces (i.e., the composition of the zones), our original, static definition of strategy-proofness does not explicitly cover this complexity. Nevertheless, the informal argument used in the proof of Theorem 8 presented in Appendix A remains sound: no matter how other players behave over time, an officer cannot secure a better outcome by misreporting, so truth-telling continues to be a weakly dominant strategy.
@@
It is also worth noting that the efficiency gains achieved with the Dynamic Modular Priority Mechanism come at a time cost. Whereas the standard procedure requires only a single round of simultaneous preference elicitation, the dynamic approach proceeds sequentially, allocating officers one by one. While this added complexity may be negligible in small markets, it can become impractical in larger settings, where the number of officers is substantial.
@@ 7 Conclusion
We explored how designing mechanisms that restrict the preferences participants can report helps reconcile multiple policy objectives—particularly distributional constraints—with fairness principles grounded in strict priority orders. We introduced the concept of visible fairness, in which a mechanism never produces an allocation that appears to violate a participant’s priority based on the partially observed preferences. Theorem 1 demonstrated that every visibly fair mechanism operates as a variant of serial dictatorship adapted to partial preferences. We further showed that this framework can accommodate diverse distributional objectives by employing modular upper-bounds, which limit how many individuals of certain types can be placed in specified subsets of positions. Central to this approach are modular-induced message spaces, which prevent participants from specifying cross-group preference comparisons that would otherwise undermine these quotas.
We then characterized what makes these mechanisms incentive-compatible, by parsing out two critical conditions—expressiveness and (weak) availability—that together ensure no participant can profit by misreporting her partial preferences. While restricting the scope of preference reporting can, in principle, worsen efficiency, we introduced a dynamic modular-priority framework that sequentially elicits partial preferences and updates feasibility constraints in real time, thus recovering the best possible Pareto outcomes subject to capacity and distributional limits. Overall, this paper highlights how deliberately constraining participants’ message spaces provides policy makers with new levers for achieving desired policy goals while preserving fairness and incentive properties in priority-based allocation settings
@@ A.1 Proof of Theorem 1
A mechanism $\psi$ is visibly fair if and only if it is a $m$-queue allocation mechanism.
@@
Part 1: A $m$-queue allocation mechanism $\psi$ is visibly fair.
Case 1. Consider any $m\in M$ and suppose by contradiction that for individual $i$ there exists $\psi(m)_{j}\succ_{i}\psi(m)_{i}$, where $\pi(i)<\pi(j)$.
By construction of the mechanism, suppose individual $i$ is assigned at step $k$; note that both $\psi(m)_{j}\in S^{k}$ and $\psi(m)_{i}\in S^{k}$ as $\pi(i)<\pi(j)$. It follows, by $\psi(m)_{j}\succ_{i}\psi(m)_{i}$ that $\psi(m)_{i}\not\in G(S^{k},m_{k})$ — a contradiction.
Case 2. Consider any $m\in M$ and suppose by contradiction that for individual $i$ there exists $s\succ_{i}\psi(m)_{i}$, where $|\{i\in I:\psi(m)_{i}=s\}|<q_{s}$.
By construction of the mechanism, suppose individual $i$ is assigned at step $k$; note that both $s\in S^{k}$ and $\psi(m)_{i}\in S^{k}$ as $|\{i\in I:\psi(m)_{i}=s\}|<q_{s}$. It follows, by $s\succ_{i}\psi(m)_{i}$ that $\psi(m)_{i}\not\in G(S^{k},m_{k})$ — a contradiction.
@@
Take an arbitrary visibly fair mechanism $\psi$ and fix a message profile $m\in M$. Denote the resulting allocation by $a=\psi(m)$. We now construct for this profile the $m$-queue allocation procedure that replicates $a$.
Define set $S^{k}$ for each officer $k\in I$ as follows:
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): Set $s^{k}=a_{k}$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
Notice that for officer $k$, the set $S^{k}$ consists of only those states that still have remaining capacities to accommodate $k$ after the allocation of higher-ranking officers is taken into account. That is, each state $s\in S\setminus S^{k}$ is filled by higher ranking offers, $|\{i\in I:i<k\text{ and }a_{i}=s\}|=q_{s}$. Now we must have that $a^{k}\in G(S^{k},m_{k})$, otherwise there exists $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$, which makes the outcome visibly unfair as a higher ranking offer does not occupy this state.
An $m$-queue mechanism working down the priority list, using sets $\{S^{k}\}_{k=1}^{n}$ defined above and the procedure given in Definition 3, would give every officer the same assignment as $\psi$. Since the construction can be repeated for every profile $m$, $\psi$ is an $m$-queue allocation mechanism. ∎
@@ A.2 Proof of Theorem 2
The proof follows the same logic as in Theorem 1, Part 1, and is omitted for brevity.
@@
Fix an arbitrary message profile $m\in M$ and let $a=\psi(m)$ be the allocation produced by any visibly fair mechanism $\psi$. We construct zone selection functions $\{C_{k}\}_{k=1}^{n}$ and sets $\{S^{k}\}_{k=1}^{n}$ that would reproduce the same allocation step by step.
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): If $a_{k}\in z$, set $\mathcal{C}_{k}(S^{k},m)=z$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
After performing the construction for every $k$ we have defined all zone selection functions $\{C_{k}\}_{k=1}^{n}$ and sets $\{S^{k}\}_{k=1}^{n}$.
Notice that for officer $k$, the set $S^{k}$ consists of only those states that still have remaining capacities to accommodate $k$ after the allocation of higher-ranking officers is taken into account. That is, each state $s\in S\setminus S^{k}$ is filled by higher ranking offers, $|\{i\in I:i<k\text{ and }a_{i}=s\}|=q_{s}$. Now we must have that $a^{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m)$, otherwise there exists $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$, which makes the outcome visibly unfair as a higher ranking offer does not occupy this state.
A partitioned priority mechanism working down the priority list, using zone selection functions $\{C_{k}\}_{k=1}^{n}$, sets $\{S^{k}\}_{k=1}^{n}$ defined above and the procedure given in Definition 4, would assign every officer the same post as $\psi$.
Since the construction can be repeated for every profile $m$, $\psi$ is a partitioned priority mechanism. ∎
@@ A.3 Proof of Theorem 3
The proof follows the same logic as in Theorem 1, Part 1, and is omitted for brevity.
@@
Theorem 2’s constructed zone selection function and argument apply here as well. Fix an arbitrary message profile $m\in M$ and let $a=\psi(m)$ be the allocation produced by any visibly fair mechanism $\psi$. We construct zone selection functions $\{C_{k}\}_{k=1}^{n}$ and sets $\{S^{k}\}_{k=1}^{n}$ that would reproduce the same allocation step by step.
  • Step 0: Set $S^{1}=S$.
  • Step $k$ ($1\leq k\leq n$): If $a_{k}\in z$, set $\mathcal{C}_{k}(S^{k},m)=z$. If the number of officers assigned to $s^{k}$ reaches $q_{s^{k}}$, that is $\left|\{i\in I:i\leq k\text{ and }a_{i}=s^{k}\}\right|=q_{s^{k}}$, then $S^{k+1}\equiv S^{k}\backslash\{s^{k}\}$. Otherwise, $S^{k+1}=S^{k}$.
After performing the construction for every $k$ we have defined all zone selection functions $\{C_{k}\}_{k=1}^{n}$ and sets $\{S^{k}\}_{k=1}^{n}$.
Notice that for officer $k$, the set $S^{k}$ consists of only those states that still have remaining capacities to accommodate $k$ after the allocation of higher-ranking officers is taken into account. That is, each state $s\in S\setminus S^{k}$ is filled by higher ranking offers, $|\{i\in I:i<k\text{ and }a_{i}=s\}|=q_{s}$. Now we must have that $a^{k}\in G(S^{k},m_{k})\cap\mathcal{C}_{k}(S^{k},m)$, otherwise there exists $s\in S^{k}$ such that $s\succ_{m_{k}}a_{k}$, which makes the outcome visibly unfair as a higher ranking offer does not occupy this state.
Also, see that the constructed zone selection function is a ranked zone selection function. Suppose not, then for officer $k\in I$, $S^{k}\cap\mathcal{C}_{k}(S^{k},m)=\left\{min\left(\mathcal{C}_{k}(S^{k},m),m_{k}\right)\right\}$ and there is a zone $z\in Z$ such that $z\rhd_{k}\mathcal{C}_{k}(S^{k},m)$ and $max\left(m_{k},z\right)\in S^{k}$. Since $max\left(m_{k},z\right)\succ_{m_{k}}min\left(\mathcal{C}_{k}(S^{k},m),m_{k}\right)$, $a^{k}=min\left(\mathcal{C}_{k}(S^{k},m_{k}),m_{k}\right)$ is visibly unfair for officer $k$ under $m$.
A ranked partitioned priority mechanism working down the priority list, using ranked zone selection functions $\{C_{k}\}_{k=1}^{n}$, sets $\{S^{k}\}_{k=1}^{n}$ defined above and the procedure given in Definition 5, would assign every officer the same post as $\psi$.
Since the construction can be repeated for every profile $m$, $\psi$ is a ranked partitioned priority mechanism.